Find the hcfof 210 and 55 and express as a linear combination

Grade:10

1 Answers

Arun
6 years ago
HCF of 210 & 55
210 = 55* 3 + 45   ….....(i)
55 = 45 * 1 +10  ….........(ii)
45 = 10 *4 +5  …...........(iii)
10 = 5 *2 + 0
hence HCF of 210 & 55 = 5
 
now from (iii), we get
4 5 = 10 *4 + 5
so 
5 = 45 – 10*4
5 = 45 – (55 – 45)*4
5 = 45 – 55*4 + 45*4
5 = 45 *5 – 55*4
5 = (210 – 55*3) *5 – 55*4
5 = 210*5 – 55*15 – 55*4
5 = 210*5 – 55*19
5 = 210 x + 55 y
where x = 5, y = –19